# Parameter-Efficient Subspace Optimization for LLM Fine-Tuning

### v1

@@ Abstract
This paper develops a new perspective on parameter-efficient fine-tuning for LLMs, inspired by the classical theory of subspace minimization. We introduce a unifying framework, Parameter-Efficient Subspace Optimization (PESO), which not only recovers many existing methods such as LoRA but also bridges them with the principled algorithmic and theoretical foundations of subspace optimization. This connection highlights a natural “exploration–exploitation” view of subspace methods, guiding the design of new algorithms that achieve strong convergence performance while still preserving memory efficiency. Importantly, our framework establishes the convergence in the full-parameter space, resolving a critical gap of LoRA variants where low-rank updates lack such guarantees. We further instantiate the framework into a practical algorithm named PESO-LoRA, based on LoRA-type parameterization. Our algorithm achieves notable improvements over existing methods on standard benchmarks.
@@ 1 Introduction
Pre-training and fine-tuning deep neural networks are the cornerstones of modern AI, powering the success of large-scale foundation models such as Large Language Models (LLMs) [brown2020language]. At their core, both procedures reduce to solving a high-dimensional optimization problem over weight matrices:
@@
where $\ell(\cdot)$ is the loss function, $W_{0}$ is the initialization, and $\Delta W$ the increment. In practice, (1) is typically solved by first-order methods such as Adam [kingma2014adam] and AdamW [loshchilov2017decoupled], which are the workhorses of large-scale training. However, these methods require storing additional optimizer states (e.g., momentum and velocity), and for LLMs this overhead places enormous pressure on memory resources, making parameter-efficient strategies appealing.
In the realm of fine-tuning, we often have limited labeled data for a downstream task but still wish to adapt the pretrained weights effectively and efficiently. Therefore, updating the entire parameter set is memory-intensive. This motivates the study of Parameter-Efficient Fine-Tuning (PEFT) methods [han2024parameter, houlsby2019parameter, hu2022lora], where optimization is restricted to a smaller set of parameters initialized from pretrained weights. In other words, $W_{0}$ denotes weights obtained from a large-scale pretraining phase, and $\Delta W$ is not updated freely but instead follows an efficient parameterization that constrains the search space.
A popular PEFT method is low-rank adaptation (LoRA, [hu2022lora]), where matrices in $\Delta W$ are expressed as the product of two low-rank factors. LoRA has shown strong empirical success, reducing memory costs while achieving competitive downstream performance. However, it suffers from two key limitations: 1) performance often lags behind full-parameter fine-tuning (Figure 1, left: MetaMathQA); 2) theoretical guarantees are limited, with convergence typically shown only for the low-rank factors (Figure 1, middle: a synthetic example illustrating LoRA’s potentially unbounded loss gap). To address these issues, many LoRA variants [hayou2024lora+, wang2024loraga, wang2024lora, zhang2023adalora, zhang2025one] have been proposed, yet they largely inherit the same shortcomings and leave the following fundamental question open:
@@
Figure 1: Comparison of full-parameter tuning, LoRA, and our method (PESO-LoRA). Left: MetaMathQA. Middle: synthetic example $\min_{W}\|W-M\|_{F}^{2}$ with $M=10\cdot\operatorname{diag}(1,\dots,1,0,\dots,0)$ (r+1r{+}1 ones); see Appendix A. Right: optimization trajectories. PESO-LoRA bridges the loss gap of LoRA while preserving memory and computation efficiency.
To address this question, we reveal an inherent connection between PEFT and the classical idea of subspace minimization, a long-standing nonlinear optimization strategy dating back to [conn1994iterated, cragg1969study]. The central philosophy is to decompose a large-scale problem like (1) into iterative, simpler subproblems constrained to carefully chosen subspaces. This view resonates naturally with modern PEFT methods, which restrict updates to structured low-rank forms for better efficiency. Interestingly, subspace minimization historically received less attention in the optimization society, since full-parameter information were often affordable in traditional applications. However, it is especially well suited to LLM training, where massive dimensionality calls for memory-efficient methods.
Formally, we build on the notion of intrinsic dimensionality in LLM training [aghajanyan2020intrinsic, li2018measuring], recognized in [hu2022lora] as the origin of LoRA: there exists a dimension-lifting map $\mathcal{M}:\mathbb{R}^{d}\to\mathbb{R}^{m\times n}$, with $d\ll m\times n$, such that the optimal solution $\Delta W^{*}$ of (1) satisfies
$\Delta W^{*}\approx\mathcal{M}(\xi^{*}),\quad\xi^{*}:=\arg\min_{\xi\in\mathbb{R}^{d}}\ell(W_{0}+\mathcal{M}(\xi)).$ (2)
@@
$\Delta W^{*}\approx\sum_{k}\mathcal{M}_{k}(\xi_{k}^{*}),\quad\xi_{k}^{*}:=\arg\min_{\xi\in\mathbb{R}^{d}}\ell(W_{0}+\sum_{i=1}^{k-1}\mathcal{M}_{i}(\xi_{i}^{*})+\mathcal{M}_{k}(\xi)).$ (3)
where each image of $\mathcal{M}_{k}$ approximates a subspace and $\xi_{k}^{*}$ is its low-dimensional coordinate. In essence, the complexity of $\mathcal{M}$ is captured by a sequence of piecewise-linear subspaces. This philosophy parallels classical approximation schemes in numerical analysis such as finite element methods [bathe2006finite].
@@
1.1 Contributions
Our contributions can be summarized at three levels. Although our focus is on PEFT, many of the ideas developed here naturally extend to pre-training.
@@ 1.1 Contributions
We introduce a novel framework PESO for memory-efficient training inspired by classical subspace minimization [conn1994iterated], unifying existing PEFT approaches such as LoRA variants [hu2022lora, wang2024loraga, wang2024lora, zhang2023adalora, zhang2025one] and GaLore [zhao2024galore]. This framework allows us to explore the rich algorithmic techniques in subspace methods, providing systematic guidance to improve memory-efficient methods. In particular, we highlight two complementary directions: exploration of new subspaces through information from the full gradient, and exploitation of the current subspace via streaming SVD representations.
@@
Our exploration mechanism, full gradient restart, enables the framework to effectively guide training dynamics. The resulting algorithm is, to our knowledge, the first method for LLM fine-tuning that combines the practical effectiveness of LoRA-style designs with a provable convergence to full-parameter optimality up to small errors, without requiring additional assumptions such as explicit low-rankness of the solution.
@@ 1.2 Related Work
LoRA [hu2022lora] is perhaps the most widely known PEFT method, and numerous variants have been proposed for better performance. For instance, LoRA+ [hayou2024lora+] introduces imbalanced learning rates; PiSSA [meng2024pissa] proposes an initialization from SVD of $W_{0}$; and AdaLoRA [zhang2023adalora] maintains an adaptive SVD-based low-rank representation. Other extensions focus on gradient scaling [tastan2025loft, zhang2024riemannian]. More recent work leverages information from the full gradient: LoRA-GA [wang2024loraga] and LoRA-Pro [wang2024lora] propose memory-efficient gradient approximations, and LoRA-One [zhang2025one] employs the SVD of the full gradient for initialization.
Convergence guarantees for PEFT algorithms remain scarce, and existing results typically address only the low-dimensional parameters [jiang2024unified]. A related line of work studies subspace descent methods [chen2025memory, he2024subspace, kozak2019stochastic, liang2024memory], which constrain updates to $W_{k+1}\leftarrow W_{k}-\eta_{k}P_{k}P_{k}^{\top}G_{k}$, where $P_{k}$ is the projection matrix, $\eta_{k}$ the learning rate, and $G_{k}$ the full gradient. These approaches establish convergence in the full-parameter space, but under extra structural assumptions. For example, [liang2024memory] analyze a continuous-time variant via Lyapunov arguments, but require $P^{\top}G=0\Rightarrow G=0$, which holds only if $P$ has full column rank—an unrealistic condition when $r<m$. Likewise, [chen2025memory, he2024subspace, kozak2019stochastic] rely on random projection theory, assuming $\mathbb{E}[PP^{\top}]=I_{m}$ and $P^{\top}P=I_{r}$, conditions not needed in our analysis. LDAdam [robert2024ldadam] imposes a design assumption that the projection $P$ must produce a strict contraction. Closer to LoRA, [jang2024lora] provide a convergence analysis, but only within the Neural Tangent Kernel (NTK) regime, limiting its applicability.
@@
Subspace minimization is a classical theme in nonlinear optimization [conn1994iterated, cragg1969study, yuan2014review]. It was historically overshadowed by full-parameter algorithms such as L-BFGS [liu1989limited] and conjugate gradient methods [nocedal1999numerical, Ch. 5], since many traditional applications could afford storing full gradients and quasi-Newton pairs. More recently, however, subspace-based strategies have re-emerged in large-scale derivative-free optimization [cartis2023scalable, dzahini2024stochastic, menickelly2024augmenting, nozawa2025zeroth, zhang2025scalable], where gradients are unavailable and low-dimensional surrogates are crucial. Furthermore, in the pre-training regime, several recent works have leveraged subspace techniques through projection-based methods, offering insights into effective subspace selection for LLM training. GaLore [zhao2024galore] constructs the projection subspace using a top-$r$ SVD of the full gradient, while Fira [chen2024fira] augments this with a gradient-correction step to reduce projection bias. SARA [zhang2025breaking] employs importance sampling to choose subspaces based on the SVD spectrum of the full gradient. APOLLO [zhu2024apollo] instead uses randomized projections to define the subspace, and SubTrack++ [rajabi2025subtrack++] explores Grassmannian subspaces for improved tracking.
@@ 2 PESO: A Framework From Subspace Minimization
Under representation (4), the evolution of $W_{k}$ can be viewed as three complementary operations: 1) exploration: updating $\mathcal{M}_{k}$ to select a new subspace $\mathcal{S}_{k}$, (line 6 of Algorithm 1) 2) exploitation: optimizing $\xi_{k}$ within the current subspace (line 8-9 of Algorithm 1), and 3) updating $\widetilde{W}_{k}$ to absorb past progress into the anchored weights. These operations mirror the classical paradigm of subspace minimization [conn1994iterated], where a large-scale problem is solved by iteratively: (i) constructing a subspace based on local information such as gradients, (ii) solving a reduced subproblem within that subspace, and (iii) updating the iterate to incorporate the subspace solution.
@@
Leveraging this connection to subspace minimization, we present our generic framework PESO in Algorithm 1. It is important to note that, by selecting corresponding parameterization of $\mathcal{M}_{k}$, UpdateSubspace, and Opt, we are able to recover a variety of existing benchmarking methods in parameter-efficient training; see representatives in Table 1. We also remark that Algorithm 1 is equivalent to the classical two-loop subspace minimization scheme [conn1994iterated], which we defer to Appendix B in Algorithm 4.
@@
1:Initialization $W_{0}\in\mathbb{R}^{m\times n}$, $\xi_{0}\in\mathbb{R}^{d}$, and $\mathcal{M}_{0}$; an algorithmic subroutine UpdateSubspace, an optimizer Opt, frequency $K$.
@@
5:  if $k-1\bmod K=0$ then $\triangleright$ Exploration to new $\mathcal{S}_{k}$
6:   $(\mathcal{M}_{k},\widetilde{W}_{k})\leftarrow{\texttt{UpdateSubspace}}(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
@@
8:  $\Delta\xi_{k}\leftarrow{\texttt{Opt}}(\xi_{k-1},\mathcal{M}_{k})$ $\triangleright$ Exploitation of current $\mathcal{S}_{k}$
@@ 2.1 Subspace Exploration-Exploitation in PESO
Exploring new subspaces is essential for navigating the full-parameter space under memory restriction. Algorithm 1 carries out exploration by UpdateSubspace, which updates both $\mathcal{M}_{k}$ and $\widetilde{W}_{k}$. Such updates are often performed lazily every $K$ iterations, as in prior works [liang2024memory, zhang2023adalora, zhao2024galore, zhu2024apollo].
Depending on how much $\mathcal{M}_{k}$ is changed, two philosophies arise for how exploration interacts with the low-dimensional $\xi_{k}$: warm-start and restart. These are simply two modes of UpdateSubspace:
  • Warm-start. Preserve $\xi_{k}$ and keep $\widetilde{W}_{k}$ fixed. Formally,
@@
Intuitively, warm-start provides smoother transitions when consecutive subspaces remain similar, while restart marks a new phase, useful when the optimization geometry changes sharply. In practice, these two modes naturally lead to two main approaches for designing UpdateSubspace. A warm-start typically updates the parameterization of $\mathcal{M}_{k}$ smoothly along an optimization trajectory—for example, by applying an Adam step on subspace parameters as in LoRA variants—yielding a gradually evolving subspace. Restart, on the other hand, often reassigns $\mathcal{M}_{k}$ directly using local information such as gradients. This strategy is common in classical optimization; for example, in line search (a one-dimensional subspace method) each iteration resets the step size initialization when a new direction is chosen [nocedal1999numerical, Ch. 3]. It is also used in LLM training, as in GaLore [zhao2024galore], which periodically resets the subspace via the SVD of the full gradient. Concrete examples of both approaches are summarized in Table 1, and Section 3.1 introduces a new restart scheme leveraging full gradients.
@@
Between two updates of UpdateSubspace, our framework performs $K$ iterations of Opt within the current subspace $\mathcal{S}_{k}$. This amounts to solving the subproblem
@@
The philosophy relies on a common belief in classical optimization: during an optimization procedure, once an effective subspace is identified, repeatedly exploiting it for multiple iterations improves efficiency. This principle underlies many classical optimization methods, such as trust-region methods [nocedal1999numerical, Ch. 4] and L-BFGS-B [byrd1995limited].

### v2

@@ Abstract
This paper develops a new perspective on parameter-efficient fine-tuning (PEFT) for LLMs, inspired by classical subspace minimization. We introduce a unifying framework, Parameter-Efficient Subspace Optimization (PESO), which recovers existing methods such as LoRA and connects them to the principled algorithmic and theoretical foundations of subspace optimization. This connection highlights a natural “exploration–exploitation” view of subspace methods, guiding the design of new algorithms that achieve strong convergence performance while still preserving memory efficiency. We instantiate the framework into a practical algorithm, PESO-LoRA, based on a LoRA-type parameterization. Importantly, we provide convergence guarantees stated in the full-parameter space for the induced update, addressing a key limitation of LoRA-style analyses that only track low-dimensional factors. Empirically, PESO-LoRA improves over strong PEFT baselines on standard fine-tuning benchmarks.
@@ 1 Introduction
Training deep neural networks is the cornerstone of modern AI, powering the success of large-scale foundation models such as Large Language Models (LLMs) (Brown et al., 2020). At the core, it reduces to solving a high-dimensional optimization problem over weight matrices:
@@
where $\ell(\cdot)$ is the loss function, $W_{0}$ is the initialization, and $\Delta W$ the increment. In practice, (1) is typically solved by first-order methods such as Adam (Kingma and Ba, 2014) and AdamW (Loshchilov and Hutter, 2017), which are the workhorses of large-scale training. However, these methods require storing additional optimizer states (e.g., momentum and velocity), and for LLMs this overhead places enormous pressure on memory resources, making parameter-efficient strategies appealing.
In the realm of fine-tuning, we often have limited labeled data for a downstream task but still wish to adapt the pretrained weights effectively and efficiently. Therefore, updating the entire parameter set is memory-intensive. This motivates the study of Parameter-Efficient Fine-Tuning (PEFT) methods (Han et al., 2024; Houlsby et al., 2019; Hu et al., 2022), where optimization is restricted to a smaller set of parameters initialized from pretrained weights. In other words, $W_{0}$ denotes weights obtained from a large-scale pretraining phase, and $\Delta W$ is not updated freely but instead follows an efficient parameterization that constrains the search space.
Figure 1: Comparison of full-parameter tuning, LoRA, and our method (PESO-LoRA). Left: optimization trajectories. Top Right: MetaMathQA. Bottom Right: synthetic example $\min_{W}\|W-M\|_{F}^{2}$ with $M=10\cdot\operatorname{diag}(1,\dots,1,0,\dots,0)$ (r+1r{+}1 ones); see Appendix A. PESO-LoRA bridges the loss gap of LoRA while preserving memory and computation efficiency.
A popular PEFT method is low-rank adaptation (LoRA, Hu et al. (2022)), where matrices in $\Delta W$ are expressed as the product of two low-rank factors. LoRA has shown strong empirical success, reducing memory costs while achieving competitive downstream performance. However, it suffers from two key limitations: 1) performance often lags behind full-parameter fine-tuning (Figure 1, top right: MetaMathQA); 2) theoretical guarantees are limited, with convergence typically shown only for the low-rank factors (Figure 1, bottom right: a synthetic example illustrating LoRA’s potentially unbounded loss gap). To address these issues, many LoRA variants (Hayou et al., 2024; Wang et al., 2024a, b; Zhang et al., 2023, 2025b) have been proposed, yet they largely inherit the same shortcomings and leave the following fundamental question open:
@@
To address this question, we reveal an inherent connection between PEFT and the classical idea of subspace minimization, a long-standing nonlinear optimization strategy dating back to Conn et al. (1994); Cragg and Levy (1969). The central philosophy is to decompose a large-scale problem like (1) into iterative, simpler subproblems constrained to carefully chosen subspaces. This view resonates naturally with modern PEFT methods, which restrict updates to structured low-rank forms for better efficiency. Interestingly, subspace minimization historically received less attention in the optimization society, since full-parameter information were often affordable in traditional applications. However, it is especially well suited to LLM training, where massive dimensionality calls for memory-efficient methods.
Formally, we build on the notion of intrinsic dimensionality in LLM training (Aghajanyan et al., 2020; Li et al., 2018), recognized in (Hu et al., 2022) as the origin of LoRA: there exists a dimension-lifting map $\mathcal{M}:\mathbb{R}^{d}\to\mathbb{R}^{m\times n}$, with $d\ll m\times n$, such that the optimal solution $\Delta W^{*}$ of (1) satisfies
$\Delta W^{*}\approx\mathcal{M}(\xi^{*}),\,\xi^{*}:=\arg\min_{\xi\in\mathbb{R}^{d}}\ell(W_{0}+\mathcal{M}(\xi)).$ (2)
@@
$\begin{array}[]{l}\Delta W^{*}\approx\sum_{k}\mathcal{M}_{k}(\xi_{k}^{*}),\\ \xi_{k}^{*}:=\arg\min_{\xi}\ell\!\left(W_{0}+\sum_{i=1}^{k-1}\mathcal{M}_{i}(\xi_{i}^{*})+\mathcal{M}_{k}(\xi)\right).\end{array}$ (3)
Each image of $\mathcal{M}_{k}$ approximates a subspace and $\xi_{k}^{*}$ is its low-dimensional coordinate. In essence, the complexity of $\mathcal{M}$ is captured by a sequence of piecewise-linear subspaces. This philosophy parallels classical approximation schemes in numerical analysis such as finite element methods (Bathe, 2006).
@@
Contributions. Our contributions can be summarized at three levels. Although our focus is on PEFT, many of the ideas developed here naturally extend to pre-training.
@@
We introduce a novel framework PESO for memory-efficient training inspired by classical subspace minimization (Conn et al., 1994), unifying existing PEFT approaches such as LoRA variants (Hu et al., 2022; Wang et al., 2024a, b; Zhang et al., 2023, 2025b) and GaLore (Zhao et al., 2024). This framework allows us to explore the rich algorithmic techniques in subspace methods, providing systematic guidance to improve memory-efficient methods. In particular, we highlight two complementary directions: exploration of new subspaces through information from the full gradient, and exploitation of the current subspace via streaming SVD representations.
@@
Our exploration mechanism, full gradient restart, enables the framework to effectively guide training dynamics. The resulting algorithm is tailored to LLM fine-tuning and combines the empirical efficiency of LoRA-style designs with convergence guarantees stated in the full-parameter space for the induced update. This aspect is largely absent from existing LoRA-style analyses, which typically track only the low-dimensional factors.
@@
Related Work. LoRA (Hu et al., 2022) is perhaps the most widely known PEFT method, and numerous variants have been proposed for better performance. For instance, LoRA+ (Hayou et al., 2024) introduces imbalanced learning rates; PiSSA (Meng et al., 2024) proposes an initialization from SVD of $W_{0}$; and AdaLoRA (Zhang et al., 2023) maintains an adaptive SVD-based low-rank representation. Other extensions focus on gradient scaling (Tastan et al., 2025; Zhang and Pilanci, 2024). More recent work leverages information from the full gradient: LoRA-GA (Wang et al., 2024a) and LoRA-Pro (Wang et al., 2024b) propose memory-efficient gradient approximations, and LoRA-One (Zhang et al., 2025b) employs the SVD of the full gradient for initialization.
One closely related line of work studies memory-efficient training via projection or subspace updates in the pre-training regime. GaLore (Zhao et al., 2024) constructs the projection subspace using a top-$r$ SVD of the full gradient, while Fira (Chen et al., 2024) augments this with a gradient-correction step to reduce projection bias. SARA (Zhang et al., 2025a) employs importance sampling to choose subspaces based on the SVD spectrum of the full gradient. APOLLO (Zhu et al., 2024) instead uses randomized projections to define the subspace, and SubTrack++ (Rajabi et al., 2025) explores Grassmannian subspaces for improved tracking. While these pre-training methods can be adapted to PEFT, they are not designed to exploit the strong intrinsic dimensionality (2) in PEFT tasks and can underperform LoRA-style parameterizations; see Table 5. In contrast, PESO builds on classical subspace minimization (Conn et al., 1994), embedding LoRA-type structures and viewing fine-tuning as sequential subproblems that iteratively capture intrinsic dimensionality (3).
Convergence guarantees for PEFT algorithms remain scarce, and existing results typically address only the low-dimensional parameters (Jiang et al., 2024). A closely related line of work studies randomized subspace descent (Kozak et al., 2019), including RSO (Chen et al., 2025) and GoLore (He et al., 2024), which construct subspaces via randomized projections and provide convergence guarantees in the full-parameter space. Their exploration strategies differ from ours and they typically do not encode PEFT-specific LoRA structures. Other approaches require additional structural conditions, e.g., Liang et al. (2024) assumes the projection subspace has full column rank (unrealistic when $r<m$) and LDAdam (Robert et al., 2024) assumes a strict contraction. Closer to LoRA, Jang et al. (2024) analyze LoRA only in the NTK regime, and RAC-LoRA (Malinovsky et al., 2024) proposes a randomized asymmetric-chaining variant, a different exploration mechanism from ours.
@@
Subspace minimization is a classical theme in nonlinear optimization (Conn et al., 1994; Cragg and Levy, 1969; Yuan, 2014). It was historically overshadowed by full-parameter algorithms such as L-BFGS (Liu and Nocedal, 1989) and conjugate gradient methods (Nocedal and Wright, 2006, Ch. 5), since many traditional applications could afford storing full gradients and quasi-Newton pairs. More recently, however, subspace-based strategies have re-emerged in large-scale derivative-free optimization (Cartis and Roberts, 2023; Dzahini and Wild, 2024; Menickelly, 2024; Nozawa et al., 2025; Zhang, 2025), where gradients are unavailable and low-dimensional surrogates are crucial.
@@ 2 PESO: A Framework From Subspace Minimization
Under representation (4), the evolution of $W_{k}$ can be viewed as three complementary operations: 1) exploration: updating $\mathcal{M}_{k}$ to select a new subspace $\mathcal{S}_{k}$, (line 6 of Algorithm 1) 2) exploitation: optimizing $\xi_{k}$ within the current subspace (line 8-9 of Algorithm 1), and 3) updating $\widetilde{W}_{k}$ to absorb past progress into the anchored weights. These operations mirror the classical paradigm of subspace minimization (Conn et al., 1994), where a large-scale problem is solved by iteratively: (i) constructing a subspace based on local information such as gradients, (ii) solving a reduced subproblem within that subspace, (iii) updating the iterate to incorporate the subspace solution.
@@
Leveraging this connection to subspace minimization, we present our generic framework PESO in Algorithm 1. It is important to note that, by selecting corresponding parameterization of $\mathcal{M}_{k}$, UpdateM, and Opt, we are able to recover a variety of existing benchmarking methods in parameter-efficient training; see representatives in Table 1. We also remark that Algorithm 1 is equivalent to the classical two-loop subspace minimization scheme (Conn et al., 1994), which we defer to Appendix B in Algorithm 4.
@@
1:Initialization $W_{0}\in\mathbb{R}^{m\times n}$, $\xi_{0}\in\mathbb{R}^{d}$, and $\mathcal{M}_{0}$; an algorithmic subroutine UpdateM, an optimizer Opt, frequency $K$.
@@
5:  if $k-1\bmod K=0$ then $\triangleright$ Explore new $\mathcal{S}_{k}$
6:   $(\mathcal{M}_{k},\widetilde{W}_{k})$ $\leftarrow{\texttt{UpdateM}}(\mathcal{M}_{k-1},\widetilde{W}_{k-1})$.
@@
8:  $\Delta\xi_{k}\leftarrow{\texttt{Opt}}(\xi_{k-1},\mathcal{M}_{k})$ $\triangleright$ Exploit current $\mathcal{S}_{k}$
@@ 2.1 Subspace Exploration-Exploitation in PESO
Exploring new subspaces is essential for navigating the full-parameter space under memory restriction. Algorithm 1 carries out exploration by UpdateM, which updates both $\mathcal{M}_{k}$ and $\widetilde{W}_{k}$. Such updates are often performed lazily every $K$ iterations, as in prior works (Liang et al., 2024; Zhang et al., 2023; Zhao et al., 2024; Zhu et al., 2024).
Depending on how much $\mathcal{M}_{k}$ is changed, two philosophies arise for how exploration interacts with the low-dimensional $\xi_{k}$: warm-start and restart. These are simply two modes of UpdateM:
  • Warm-start. Preserve $\xi_{k}$ and keep $\widetilde{W}_{k}$ fixed:
@@
Intuitively, warm-start provides smoother transitions when consecutive subspaces remain similar, while restart marks a new phase, useful when the optimization geometry changes sharply. In practice, these two modes naturally lead to two main approaches for designing UpdateM. A warm-start typically updates the parameterization of $\mathcal{M}_{k}$ smoothly along an optimization trajectory—for example, by applying an Adam step on subspace parameters as in LoRA variants—yielding a gradually evolving subspace. Restart, on the other hand, often reassigns $\mathcal{M}_{k}$ directly using local information such as gradients. This strategy is common in classical optimization; for example, in line search (a one-dimensional subspace method) each iteration resets the step size initialization when a new direction is chosen (Nocedal and Wright, 2006, Ch. 3). It is also used in LLM training, as in GaLore (Zhao et al., 2024), which periodically resets the subspace via the SVD of the full gradient. Concrete examples of both approaches are summarized in Table 1, and Section 3.1 introduces a new restart scheme leveraging full gradients.
@@
Between two updates of UpdateM, our framework performs $K$ iterations of Opt within the current subspace $\mathcal{S}_{k}$. This amounts to solving the subproblem
@@
The philosophy relies on a common belief in classical optimization: during an optimization procedure, once an effective subspace is identified, repeatedly exploiting it for multiple iterations improves efficiency. This principle underlies many classical optimization methods, such as trust-region methods (Nocedal and Wright, 2006, Ch. 4) and L-BFGS-B (Byrd et al., 1995).